OpenStudy (anonymous):

What is the slope of the line that passes through the points (–1, 13) and (4, –2)?

OpenStudy (lucaz):

slope is change in y over change in x slope = (13-(-2)) / (-1-4) the order doesnt matter, it could be slope = (-2-13) / (4-(-1)) slope = -15/5 or 15/-5 = -3

OpenStudy (anonymous):

thanks! @lucaz